# Nonlinear Causal Discovery through a Sequential Edge Orientation Approach

### v1

@@ 6.4 Intermediate Results at Individual Stages
Having presented the performance of our algorithm against competing methods, we now closely analyze its accuracy after each of the following four stages in Algorithm 2: (1) initial learning to learn the initial PDAG, (2) edge orientation to determine the causal direction of undirected edges in the PDAG, (3) edge deletion to remove superfluous edges, and (4) graph refinement to convert the PDAG to a DAG, if applicable, by applying the edge orientation step again. Figure 9 shows the F1 score computed after each of these four stages. As mentioned in Section 3, while our framework produces a PDAG in general, the implementation incorporates a fourth step to produce a DAG given the non-linear ANM assumption. Under the linear case, the F1 score is expected to remain unchanged after the orientation step, since the undirected edges in a CPDAG should not be oriented. The algorithm’s capabilities are best demonstrated in learning nonlinear DAGs, where the F1 score improves by 7.7% to 53.3% from the initial step to after the deletion step under nonlinear functions.
@@
It is imperative to recall that the initial graph is dense because it incorporates an extra candidate edge set ($U_{\alpha_{2}}$ on Line 2 in Algorithm 2), which may include some false positives. Although the deletion step appears to yield the greatest increase in the F1 score, the edge orientation step actually first uncovers more true positives from the undirected edges (Figure 13 in the Appendix). A detailed analysis shows that number of true positives increases by 3.3% to 23.6% after applying the orientation procedure. The extra candidate edges may contain true positive edges or edges in the true DAG that are crucial to correctly orienting undirected edges; these edges would otherwise not be recovered in the latter stages. Since the number of true positives remains unchanged after the edge deletion step, we can conclude that the deletion step correctly removes irrelevant edges. Therefore, the inclusion of additional candidate edges is essential and beneficial to our algorithm. Moreover, the significant increase of the F1 score from the initial stage evidences that SNOE can indeed enhance conventional algorithms that learn only an equivalence class.
6.5 Application to Real Data
@@
We applied all algorithms to the flow cytometry data set underlying the well-known Sachs network (Sachs et al., 2005). The data set was collected in a study aiming to infer causal pathways amongst 11 phosphorylated proteins and phospholipids by measuring their expression levels after performing knockouts and spikings. Through a meta-analysis of both biological data and published literature, the researchers derived a causal DAG that illustrates 17 causal relations among the 11 molecules. Given the high potential and broad applicability of causal learning in biological sciences, the Sachs network is one of the few verified causal graphs and is a popular means to benchmark causal learning methods. While the original data contains 7466 single-cell samples, we applied the algorithms to only the continuous version of the observational data set, which reduces the final data set to 2603 samples. It should be noted that the underlying skeleton is not fully connected; the graph consists of two disjoint clusters, one containing 8 nodes and the other containing 3.
@@
A breakdown of the performances in Table 3 shows the SNOE cross-validation approach produces a DAG closest to the ground truth, with a F1 score more than 30% higher than those of competing methods. Although its learned DAG is sparser than the true DAG, SNOE-CV has the highest F1 score and ties with SNOE-SS and CAM for the number of true positives captured. Despite missing several edges, it predicted very few false positives and thus has a lower SHD. The sample-splitting approach ranks second and differs from the cross-validation approach by predicting just one more false positive edge. NOTEARS and DAGMA were applied to both the original and standardized data, with the better results reported. Nevertheless, both methods still suffer from relatively high numbers of false negatives. Although CAM and SCORE have the densest DAGs predicted, they also have the highest counts of false positives and SHD.
@@ 7 Discussion
In this work, we present a novel algorithm for learning nonlinear causal DAGs through a sequential edge orientation framework. Specifically, we demonstrate that the edge orientation algorithm can learn the true DAG from the CPDAG by sequentially orienting the undirected edges. The framework is established on the pairwise additive noise model, a criterion we developed to ensure accurate inference of the causal direction for undirected edges just from the two nodes and their identified parent sets. The sequential orientation of edges is achieved through two key components: the likelihood ratio test, which provides a definitive decision on the causal direction of an undirected edge, and the edge ranking procedure, which recursively determines edges that follow the PANM to ensure the correctness of orientations made. These two procedures effectively address two fundamental questions for constraint-based approaches regarding how to determine the causal direction of edges and how to order edges for evaluation. We also propose two approaches to the likelihood ratio test, both of which demonstrate well-controlled type I error and high statistical power. SNOE provides a much more practical, yet still precise, alternative to kernel-based and regression-based tests for detecting nonlinear relations. Compared to competing nonlinear learning methods, SNOE exhibits robustness and high precision, which can be attributed to its reduced dependence on model assumptions. It also requires far less computational time and demonstrates strong generalizability to different data functions and distributions.
Potential extensions of this framework include learning nonlinear causal relations in the presence of hidden variables— a common challenge in constraint-based algorithms. The key is to adapt the edge ranking and the likelihood ratio test to take into account latent confounders. In addition, this work can be refined and expanded in several ways. One direction would be to demonstrate the identifiability of pairwise nodes in a PDAG through an alternative model. In our work, we determine the causal direction of an undirected edge based on the identifiability of a pairwise ANM, yet there may be other models that allow for inference of the causal direction between a pair of variables, which can be incorporated in our method of sequential edge orientation. Another promising direction is to more strategically integrate our edge orientation procedure with other learning algorithms to expand its use case, particularly with those utilizing KCI tests. The kernel-based tests can more accurately recover the skeleton in theory, while the sequential edge orientation procedure can reduce runtime by its partitioned learning approach over multiple undirected components.
@@
We thank Dr. Bingling Wang for helpful discussions. This work was supported by NSF grant DMS-2305631 and used computational resources and storage capacities on the Hoffman2 Shared Cluster hosted by the UCLA Institute for Digital Research and Education’s Research Technology Group.
@@ Appendix A Restricted Additive Noise Model
Suppose each variable $X_{j}$ is generated by an ANM $X_{j}=f_{j}(\text{PA}_{j})+N_{j}$, where $f_{j}$ is an arbitrary function of its parent variables and $N_{j}$ is an additive noise. Moreover, let $ND_{j}$ denote the non-descendants of $X_{j}$ in the underlying DAG. We denote by $L(Y)$ the distribution of a random variable $Y$. In a bivariate additive noise model for variables $X_{i},X_{j}$, Hoyer et al. (2008) have proven that if the triple $(f_{j},L(X_{i}),L(N_{j}))$ satisfies the following condition, then the causal relation is identifiable.
@@
Peters et al. (2014) then utilize this result to prove the identifiability of a DAG assuming a restricted additive noise model.
@@ Appendix B Proofs
B.1 Proof of Lemma 1
Proof  Given that $[X,Y\mid\text{pa}_{\mathcal{G}}(X),\text{pa}_{\mathcal{G}}(Y)]$ follows a pairwise ANM, $(X,Y)$ follow the set of SEMs in either (i) or (ii) in the PANM definition. The true DAG $\mathcal{G}_{0}$ is identifiable if its node variables are generated by a restricted additive noise model. To prove Lemma 1, we construct a DAG $\mathcal{G}^{*}$ on $X,Y,\text{pa}_{\mathcal{G}}(X)\cup\text{pa}_{\mathcal{G}}(Y)$, with only edges $\text{pa}_{\mathcal{G}}(X)\to X$, $\text{pa}_{\mathcal{G}}(Y)\to Y$, and between $X,Y$. Then, by the above identifiability result, the DAG $\mathcal{G}^{*}$, including the edge between $X$ and $Y$, is identifiable, which gives our result.
B.2 Proof of Theorem 1
@@
Suppose the variables (nodes) of the DAG $\mathcal{G}_{0}$ follow a restricted ANM. Let $(X,Y)$ be an undirected edge in $\mathcal{G}$ on Line 1 at any stage of Algorithm 1 with the input being the CPDAG of $\mathcal{G}_{0}$. Let $Z_{1}=\text{pa}_{\mathcal{G}}(X)$ and $Z_{2}=\text{pa}_{\mathcal{G}}(Y)$. If $\text{pa}_{\mathcal{G}_{0}}(X)=Z_{1},\text{pa}_{\mathcal{G}_{0}}(Y)=Z_{2}\cup\{X\}$ or $\text{pa}_{\mathcal{G}_{0}}(Y)=Z_{2},\text{pa}_{\mathcal{G}_{0}}(X)=Z_{1}\cup\{Y\}$, then $[X,Y|Z_{1},Z_{2}]$ follows a pairwise additive noise model.
@@
Lemma 2 identifies a type of undirected edges, but not all, that satisfy the PANM. We will show that the PDAG $\mathcal{G}$ contains at least one such undirected edge on Line 1 at any iteration of the algorithm. Now we prove the theorem:
@@
Proof To prove that the sequential edge orientation procedure can recover the true DAG $\mathcal{G}_{0}$ by correctly orienting all undirected edges in the CPDAG $\mathcal{E}$, we demonstrate that the following holds true for $\mathcal{G}$ at every iteration at Line 1 in Algorithm 1: If $\mathcal{G}$ is not a DAG, then there exists an undirected edge $X-Y$ such that $[X,Y\mid Z_{1},Z_{2}]$ satisfies the PANM, where $Z_{1}=\text{pa}_{\mathcal{G}}(X)$ and $Z_{2}=\text{pa}_{\mathcal{G}}(Y)$. It is easy to see that every orientation step in Algorithm 1 will only lead to correct orientation that is consistent with the DAG $\mathcal{G}_{0}$ if the input $\mathcal{G}$ is the CPDAG.
@@
Let $T$ be an undirected component of the PDAG $\mathcal{G}$ of size $|T|\geq 2$. Let $v_{1}\in T$ be a node that precedes all other nodes in $T$ according to some topological ordering $\prec$ of $\mathcal{G}_{0}$. Because $|T|\geq 2$, the neighbor set $\text{ne}_{\mathcal{G}}(v_{1})$ is not empty. If there exists $v_{2}\in\text{ne}_{\mathcal{G}}(v_{1})$ such that $\{v_{1},v_{2}\}$ satisfies the conditions in Lemma 2, then $(v_{1},v_{2})$ satisfies the PANM and the proof is complete. By construction all parents of $v_{1}$ have been identified in $\mathcal{G}$, i.e. $\text{pa}_{\mathcal{G}}(v_{1})=\text{pa}_{\mathcal{G}_{0}}(v_{1})$. Let $v_{2}$ be a node that precedes all other nodes in $\text{ne}_{\mathcal{G}}(v_{1})$ according to some sort of $\mathcal{G}_{0}$. Then, $v_{1}\in\text{pa}_{\mathcal{G}_{0}}(v_{2})$ and none of the nodes in $\text{ne}_{\mathcal{G}}(v_{1})$ is a parent of $v_{2}$ in $\mathcal{G}_{0}$. It remains to show that $\text{pa}_{\mathcal{G}_{0}}(v_{2})=\text{pa}_{\mathcal{G}}(v_{2})\cup\{v_{1}\}$. If this is not the case, then there must exist another node $v_{j}\in\text{ne}_{\mathcal{G}}(v_{2})\setminus\{v_{1}\}$ that is a parent of $v_{2}$ in $\mathcal{G}_{0}$ and $v_{j}\not\in\text{ne}_{\mathcal{G}}(v_{1})$, i.e. there is no undirected edge between $v_{1}$ and $v_{j}$ in the PDAG $\mathcal{G}$. There are two possibilities with respect to the connectivity between $v_{1}$ and $v_{j}$ in $\mathcal{G}_{0}$. The first possibility is that there is no edge between $v_{1}$ and $v_{j}$ in $\mathcal{G}_{0}$. This would form a new v-structure $v_{1}\rightarrow v_{2}\leftarrow v_{j}$ in $\mathcal{G}_{0}$, which is a contradiction to that the input $\mathcal{G}$ was the CPDAG of $\mathcal{G}_{0}$. The second possibility is that $v_{1}\to v_{j}$ in $\mathcal{G}_{0}$ since $v_{1}\prec v_{j}\in T$ and this edge has been oriented so in $\mathcal{G}$. Since neither $v_{1}-v_{2}$ nor $v_{2}-v_{j}$ has been oriented in $\mathcal{G}$, the edge $v_{1}\to v_{j}$ must have been oriented either by Line 1 or by Meek’s rule 1 on Line 1. In what follows, we show that both scenarios would lead to contradictions, and thus such $v_{j}$ does not exist.
@@
If $v_{1}\rightarrow v_{j}$ is oriented by Line 1, then there must be another node $v_{i}$ where $v_{1}-v_{i}$ was oriented in either direction on Line 1 first and $v_{i}$ is adjacent to $v_{j}$. The algorithm would also orient $v_{i}\rightarrow v_{j}$ by Line 1 or from previous actions. Shown in Figure 10 (a) and (b), there are two cases regarding the adjacency of $v_{2}$ and $v_{i}$ in $\mathcal{G}$ under $v_{1}\rightarrow v_{i}$. (a) If $v_{i},v_{2}$ are adjacent, $v_{i}$ must be a parent of $v_{2}$ because otherwise there would be a directed cycle $v_{2}\to v_{i}\to v_{j}\to v_{2}$ in $\mathcal{G}_{0}$. Then, Line 1 would orient $v_{1}\rightarrow v_{2}$ too, contradicting to that $v_{1}-v_{2}$ is undirected in $\mathcal{G}$. (b) If $v_{i},v_{2}$ are not adjacent, then the algorithm would orient $v_{j}\rightarrow v_{2}$ by Meek’s rule 1 in the following step, contradicting to that $v_{j}\in\text{ne}_{\mathcal{G}}(v_{2})$ (i.e. $v_{j}-v_{2}$ in $\mathcal{G}$). Similar arguments under $v_{i}\rightarrow v_{1}$ result in contraditions that supposedly undirected edges in $\mathcal{G}$ would have been oriented.
@@
Figure 10: First possibility: orienting $v_{1}\rightarrow v_{j}$ by Line 1 in Algorithm 1 results in contradictions where $v_{1}\rightarrow v_{2}$ is oriented in (a) and $v_{j}\rightarrow v_{2}$ in (b).
If $v_{1}\rightarrow v_{j}$ is oriented by Meek’s rule 1 on Line 1, then there must exist a node $Z\in V$ in the configuration $Z\rightarrow v_{1}-v_{j}$ and $Z$ is not adjacent to $v_{j}$. There are four possible cases, depicted in Figure 11, with respect to the possible connection between $Z$ and $v_{2}$ in the PDAG $\mathcal{G}$. Case 1: there is no edge between $Z$ and $v_{2}$. The undirected edge $v_{1}-v_{2}$ would then be oriented by rule 1 as $v_{1}\to v_{2}$, which leads to a contradiction. Case 2: The two nodes are connected by an undirected edge $Z-v_{2}$. Then, $Z\in T$ and is a parent node of $v_{1}$, contradicting the fact that $v_{1}$ precedes all other nodes in $T$. Case 3: $Z\rightarrow v_{2}$ in $\mathcal{G}$. Meek’s rule 1 would then orient $v_{2}\rightarrow v_{j}$, which would result in an incorrect orientation as $v_{j}$ is assumed to be the parent of $v_{2}$ in $\mathcal{G}_{0}$, again a contradiction. Case 4: $v_{2}\rightarrow Z$ in $\mathcal{G}$. Then the edge $v_{1}-v_{2}$ must be $v_{2}\to v_{1}$ in $\mathcal{G}_{0}$, which contradicts the assumed ordering $v_{1}\prec v_{2}$.

### v2

@@ 6.4 Intermediate Results at Individual Stages
Having presented the performance of our algorithm against competing methods, we now closely analyze its accuracy after each of the following four stages in Algorithm 2: (1) initial learning to learn the initial PDAG, (2) edge orientation to determine the causal direction of undirected edges in the PDAG, (3) edge deletion to remove superfluous edges, and (4) graph refinement to convert the PDAG to a DAG, if applicable, by applying the edge orientation step again. As mentioned in Section 3, while our framework produces a PDAG in general, the implementation incorporates a fourth step to produce a DAG given the non-linear ANM assumption. Figure 10 shows the F1 score computed after each of these four stages. The algorithm’s capabilities are best demonstrated in learning nonlinear DAGs, where the F1 score improves by 7.7% to 53.3% from the initial step to after the deletion step under nonlinear functions.
@@
It is imperative to recall that the initial graph is dense because it incorporates an extra candidate edge set ($U_{\alpha_{2}}$ on Line 2 in Algorithm 2), which may include some false positives. Although the deletion step appears to yield the greatest increase in the F1 score, the edge orientation step actually first uncovers more true positives from the undirected edges (see Figure 15 in the Appendix). A detailed analysis shows that the number of true positives increases by 3.3% to 23.6% after applying the orientation procedure. The extra candidate edges may contain true positive edges or edges in the true DAG that are crucial to correctly orienting undirected edges; these edges would otherwise not be recovered in the latter stages. Since the number of true positives remains unchanged after the edge deletion step, we can conclude that the deletion step correctly removes irrelevant edges. Therefore, the inclusion of additional candidate edges is essential and beneficial to our algorithm. Moreover, the significant increase of the F1 score from the initial stage evidences that SNOE can indeed enhance conventional algorithms that learn only an equivalence class.
Since in practice we do not know whether the nonlinear ANM is satisfied, we include the Gaussian linear case to show the performance when the identifiability result does not hold. The ground truth for evaluating this setting is the true CPDAG. In our algorithm design, we use the PC algorithm coupled with the partial correlation test to learn the CPDAG, yielding the relatively high F1 score after the first stage in Figure 10. For several networks, the F1 score also increases after the second stage of orienting edges. In most practical applications, the estimated CPDAG is not perfect and thus may not capture all v-structures. The observed increase results from recovering missing compelled edges in the true CPDAG by the likelihood ratio test, as confirmed by the increase in true positives after stage two in these networks (see Figure 15 in Appendix C.3). Since the first stage estimates a dense graph, the third stage removes false positives and thus improves the F1 score. This is helpful for the Gaussian linear case as well. Note that the final output is a DAG because we assume the nonlinear ANM; we then see a decrease in the F1 score after stage four which orients all remaining undirected edges. This is expected for Gaussian linear DAGs. While this is expected for Gaussian linear DAGs, SNOE still exhibits high accuracy in this setting, tdemonstrating its versatility for causal learning under various functional forms.
Furthermore, we conducted an experiment to study the effect of misranking edges in the orientation procedure. We evaluated the undirected edges for orientation separately in ranked and arbitrary orders on the same initial graph, and then recorded the F1 score after the orientation stage. The results in Table 4 in Appendix C.5 confirm the advantages of defining an order in our sequential method over a random search and the robustness against misranking. The analysis reveals that the ranking procedure indeed led to more true positives and higher F1 scores in the orientation stage. When edges are misranked, orienting an edge XYX-Y not fulfilling the PANM could lead to an erroneous orientation. However, the likelihood ratio test largely mitigates the impact of misranking since it may find models $X\rightarrow Y$ and $Y\rightarrow X$ indistinguishable or comparably fitted, thus leaving the edge undirected until evaluation at a later iteration. See Appendix C.5 for a more detailed discussion.
We have also examined the impact of the initial graph on the performance of our algorithm. The detailed results can be found in Appendix C.2. Furthermore, we provide an analysis and discussion on the runtime of the intermediate steps in Appendix C.4.
7 Real Data Applications
7.1 Flow-Cytometry Data
@@
We applied all methods to the flow cytometry data set underlying the well-known Sachs network (flow_cytometry). The data set was collected in a study aiming to infer causal pathways amongst 11 phosphorylated proteins and phospholipids by measuring their expression levels after performing knockouts and spikings. Through a meta-analysis of both biological data and published literature, the researchers constructed a causal DAG that illustrates 17 causal relations among the 11 molecules. Given the high potential and broad applicability of causal learning in biological sciences, the Sachs network is one of the few verified causal graphs and is a popular means to benchmark causal learning methods. While the original data contains 7466 single-cell samples, we applied the algorithms to only the continuous version of the observational data set, which reduces the final data set to 2603 samples. It should be noted that the underlying skeleton is not fully connected; the graph consists of two disjoint clusters, one containing 8 nodes and the other containing 3.
@@
The performance summary of the estimated graphs in Table 2 shows the SNOE cross-validation approach produces a DAG closest to the ground truth, with a F1 score more than 30% higher than those of competing methods. Although its learned DAG is sparser than the true DAG, SNOE-CV has the highest F1 score and ties with SNOE-SS and CAM for the number of true positives captured. Despite missing several edges, it predicted very few false positives and thus has a lower SHD. The sample-splitting approach ranks second and differs from the cross-validation approach by predicting just one more false positive. NOTEARS and DAGMA were applied to both the original and standardized data, with the better results reported. Nevertheless, both methods still suffer from relatively high numbers of false negatives. Although CAM and SCORE have the densest DAGs predicted, they also have the highest counts of false positives and SHD.
7.2 Tübingen Cause-Effect Pairs
The Cause-Effect database is a collection of 108 different cause-effect pairs sourced from various domains such as biology, climate science, economics, and sociology (mooij2016distinguishing). Each data set consists of two variables with a known causal relation. The database has emerged as a popular benchmark for evaluating bivariate causal discovery methods given its diversity in data sources and data types. In this experiment, we applied both variants of the likelihood ratio test, the sample-splitting (SS) and cross-validation (CV) approaches, to each pair using a significance level of $\alpha=0.05$. We consider 98 data sets since nine of them contain multi-dimensional variables.
The first two columns of Table 3 show the results of applying the likelihood ratio test with $\alpha=0.05$ to determine the edge orientation. As shown in the first column, for more than 60 data sets, the causal direction cannot be determined by the likelihood ratio test at the significance level of 0.05. The causal models for $X\rightarrow Y$ and for $Y\rightarrow X$ were found to be equivalent for these data sets, suggesting that the SEMs may be approximated by linear models. Further analysis reveals that the average correlation amongst these undetermined pairs is 0.49, which signals a strong linear relation. For the other data sets, both approaches achieved a high accuracy of inferring the correct causal direction for about $70\%$ of the data sets.
Approach # of undetermined cases Accuracy when causal direction is determined Overall accuracy
LRT-$S$ 63 68.6% 54.1%
LRT-CV 66 71.9% 66.3%
Table 3: Likelihood ratio test results on 98 cause-effect data sets.
mooij2016distinguishing applied their causal discovery methods to these data sets, assuming nonlinear ANMs as well. Their methods predicted the causal relations for all 98 data sets by choosing the direction with less dependence between the residual and parent variables. To compare with their methods, we chose the direction having a larger log-likelihood value as the predicted causal direction for each data set, regardless of the test significance. We focus on their results for six entropy-based approaches, since the entropy measures are closely related to the normalized MI used in our work. Only one of their six approaches reached an accuracy of around $70\%$ and the remaining all scored 40%–60%. As reported in the last column of Table 3, while LRT-SS exhibits a similar overall accuracy of 54.1%, LRT-CV outperforms most of their approaches with an overall accuracy of 66.3%. mooij2016distinguishing attribute the large variation in accuracy to discretization effects when calculating the differential entropy. In our procedure, we normalize mutual information to avoid extreme values arising from distribution skewness or discretization. Furthermore, the MI measure is only used for checking adherence to the PANM criterion, while the likelihood ratio test determines the causal relation in a robust way.
@@ 8 Discussion
In this work, we present a novel algorithm for learning nonlinear causal DAGs through a sequential edge orientation framework. Specifically, we demonstrate that the edge orientation algorithm can learn the true DAG from the CPDAG by sequentially orienting the undirected edges. The framework is established on the pairwise additive noise model, a criterion we developed to ensure accurate inference of the causal direction for undirected edges from just the two nodes and their identified parent sets. The sequential orientation of edges is achieved through two key components: the likelihood ratio test, which provides a definitive decision on the causal direction of an undirected edge, and the edge ranking procedure, which recursively determines edges that follow the PANM to ensure the correctness of orientations made. These two procedures effectively address two fundamental questions for constraint-based approaches regarding how to determine the causal direction of edges and how to order edges for evaluation. We also propose two approaches to the likelihood ratio test, both of which demonstrate well-controlled type I error and high statistical power. SNOE provides a practical, yet still precise, alternative to kernel-based and regression-based learning of nonlinear causal relations. Compared to competing methods, SNOE exhibits robustness and high precision, which can be attributed to its reduced dependence on model assumptions. It also requires far less computational time and demonstrates strong generalizability to different data functions and distributions.
Potential extensions of this framework include learning nonlinear causal relations in the presence of hidden variables— a common challenge in constraint-based algorithms. The key is to adapt the edge ranking and the likelihood ratio test to take into account latent confounders. In addition, this work can be refined and expanded in several ways. One direction is to expand our algorithm to the post-nonlinear (PNL) model, where the SEM is given as $X_{i}=g_{i}(f_{i}(\text{PA}_{i})+\varepsilon_{i})$ and the causal direction can be determined by testing the independence between $\varepsilon_{i}$ and $\text{PA}_{i}$ (zhang2009pnl_identifiability). We can adapt the pairwise ANM criterion to the PNL model, as it also relies on the independence noise property for identification, and thereby produce an evaluation order of edges in learning the true DAG from its MEC. Deep learning methods (uemura2020estimation_pnl) have been developed to better approximate $f_{i},g_{i}$ and the noise term, which would allow us to accurately compute the normalized MI, rank edges, and perform the likelihood ratio test under the PNL model. Another promising direction is to integrate alternative independence measures and estimation methods into our algorithm. For general nonlinear models, a kernel-based test (kci_test) may be more accurate and an alternative likelihood estimation, such as the method proposed in khemakhem2021causal, can be used in place of the likelihood ratio test for determining the causal direction.
@@
We thank Dr. Bingling Wang for helpful discussions. This work was supported by NSF grant DMS-2305631 and NIH grant R01GM163245. It used computational resources and storage capacities on the Hoffman2 Shared Cluster hosted by the UCLA Institute for Digital Research and Education’s Research Technology Group.
@@ Appendix A Restricted Additive Noise Model
Suppose each variable $X_{j}$ is generated by an ANM $X_{j}=f_{j}(\text{PA}_{j})+N_{j}$, where $f_{j}$ is an arbitrary function of parent variables $\text{PA}_{j}$ and $N_{j}$ is an additive noise. Moreover, let $ND_{j}$ denote the non-descendants of $X_{j}$ in the underlying DAG. We denote by $L(Y)$ the distribution of a random variable $Y$. In a bivariate additive noise model for variables $X_{i},X_{j}$, hoyer2008nonlinear have proven that if the triple $(f_{j},L(X_{i}),L(N_{j}))$ satisfies the following condition, then the causal relation is identifiable.
@@
resit_anm then utilize this result to prove the identifiability of a DAG assuming a restricted additive noise model.
@@ Appendix B Proofs
B.1 Proof of Theorem 1
@@
Suppose the variables (nodes) of the DAG $\mathcal{G}_{0}$ follow a restricted ANM. Let $(X,Y)$ be an undirected edge in $\mathcal{G}$ on Line 1 at any stage of Algorithm 1 with the input being the CPDAG of $\mathcal{G}_{0}$. Let $Z_{1}=\text{pa}_{\mathcal{G}}(X)$ and $Z_{2}=\text{pa}_{\mathcal{G}}(Y)$. If $\text{pa}_{\mathcal{G}_{0}}(X)=Z_{1},\text{pa}_{\mathcal{G}_{0}}(Y)=Z_{2}\cup\{X\}$ or $\text{pa}_{\mathcal{G}_{0}}(Y)=Z_{2},\text{pa}_{\mathcal{G}_{0}}(X)=Z_{1}\cup\{Y\}$, then $X,Y|Z_{1},Z_{2}]$ follows a pairwise additive noise model.
@@
Lemma 2 identifies a type of undirected edge, amongst all, that satisfy the PANM. We will show that the PDAG $\mathcal{G}$ contains at least one such undirected edge on Line 1 at any iteration of the algorithm. Now we prove the theorem:
@@
Proof To prove that the sequential edge orientation procedure can recover the true DAG $\mathcal{G}_{0}$ by correctly orienting all undirected edges in the CPDAG $\mathcal{E}$, we demonstrate that the following holds true for $\mathcal{G}$ at every iteration at Line 1 in Algorithm 1: If $\mathcal{G}$ is not a DAG, then there exists an undirected edge XYX-Y such that $X,Y\mid Z_{1},Z_{2}]$ satisfies the PANM, where $Z_{1}=\text{pa}_{\mathcal{G}}(X)$ and $Z_{2}=\text{pa}_{\mathcal{G}}(Y)$. It is easy to see that every orientation step in Algorithm 1 will only lead to correct orientation that is consistent with the DAG $\mathcal{G}_{0}$ if the input $\mathcal{G}$ is the CPDAG.
@@
Let $T$ be an undirected component of the PDAG $\mathcal{G}$ of size $|T|\geq 2$. Let $v_{1}\in T$ be a node that precedes all other nodes in $T$ according to some topological ordering $\prec$ of $\mathcal{G}_{0}$. Because $|T|\geq 2$, the neighbor set $\text{ne}_{\mathcal{G}}(v_{1})$ is not empty. If there exists $v_{2}\in\text{ne}_{\mathcal{G}}(v_{1})$ such that $\{v_{1},v_{2}\}$ satisfies the conditions in Lemma 2, then $(v_{1},v_{2})$ satisfies the PANM and the proof is complete. By construction all parents of $v_{1}$ have been identified in $\mathcal{G}$, i.e. $\text{pa}_{\mathcal{G}}(v_{1})=\text{pa}_{\mathcal{G}_{0}}(v_{1})$. Let $v_{2}$ be a node that precedes all other nodes in $\text{ne}_{\mathcal{G}}(v_{1})$ according to some sort of $\mathcal{G}_{0}$. Then, $v_{1}\in\text{pa}_{\mathcal{G}_{0}}(v_{2})$ and none of the nodes in $\text{ne}_{\mathcal{G}}(v_{1})$ is a parent of $v_{2}$ in $\mathcal{G}_{0}$. It remains to show that $\text{pa}_{\mathcal{G}_{0}}(v_{2})=\text{pa}_{\mathcal{G}}(v_{2})\cup\{v_{1}\}$. If this is not the case, then there must exist another node $v_{j}\in\text{ne}_{\mathcal{G}}(v_{2})\setminus\{v_{1}\}$ that is a parent of $v_{2}$ in $\mathcal{G}_{0}$ and $v_{j}\not\in\text{ne}_{\mathcal{G}}(v_{1})$, i.e. there is no undirected edge between $v_{1}$ and $v_{j}$ in the PDAG $\mathcal{G}$. There are two possibilities with respect to the connectivity between $v_{1}$ and $v_{j}$ in $\mathcal{G}_{0}$. The first possibility is that there is no edge between $v_{1}$ and $v_{j}$ in $\mathcal{G}_{0}$. This would form a new v-structure $v_{1}\rightarrow v_{2}\leftarrow v_{j}$ in $\mathcal{G}_{0}$, which is a contradiction to that the input $\mathcal{G}$ is the CPDAG of $\mathcal{G}_{0}$. The second possibility is that $v_{1}\to v_{j}$ in $\mathcal{G}_{0}$ since $v_{1}\prec v_{j}\in T$ and this edge has been oriented so in $\mathcal{G}$. Since neither $v_{1}-v_{2}$ nor $v_{2}-v_{j}$ has been oriented in $\mathcal{G}$, the edge $v_{1}\to v_{j}$ must have been oriented either by Line 1 or by Meek’s rule 1 on Line 1. In what follows, we show that both scenarios would lead to contradictions, and thus such $v_{j}$ does not exist.
@@
If $v_{1}\rightarrow v_{j}$ is oriented by Line 1, then there must be another node $v_{i}$ such that $v_{1}-v_{i}$ was oriented in either direction on Line 1 first and $v_{i}$ is adjacent to $v_{j}$. The algorithm would also orient $v_{i}\rightarrow v_{j}$ by Line 1 or from previous actions. Shown in Figure 11 (a) and (b), there are two cases regarding the adjacency of $v_{2}$ and $v_{i}$ in $\mathcal{G}$ assuming $v_{1}\rightarrow v_{i}$ has been oriented. (a) If $v_{i},v_{2}$ are adjacent, $v_{i}$ must be a parent of $v_{2}$ because otherwise there would be a directed cycle $v_{2}\to v_{i}\to v_{j}\to v_{2}$ in $\mathcal{G}_{0}$. Then, Line 1 would orient $v_{1}\rightarrow v_{2}$ too, contradicting to that $v_{1}-v_{2}$ is undirected in $\mathcal{G}$. (b) If $v_{i},v_{2}$ are not adjacent, then the algorithm would orient $v_{j}\rightarrow v_{2}$ by Meek’s rule 1 in the following step, contradicting to that $v_{j}\in\text{ne}_{\mathcal{G}}(v_{2})$ (i.e. $v_{j}-v_{2}$ in $\mathcal{G}$). Similar arguments under the orientation $v_{i}\rightarrow v_{1}$ result in contradictions that supposedly undirected edges in $\mathcal{G}$ would have been oriented.
If $v_{1}\rightarrow v_{j}$ is oriented by Meek’s rule 1 on Line 1, then there must exist a node $Z\in V$ in the configuration $Z\rightarrow v_{1}-v_{j}$ and $Z$ is not adjacent to $v_{j}$. There are four possible cases, depicted in Figure 12, with respect to the connection between $Z$ and $v_{2}$ in the PDAG $\mathcal{G}$. Case 1: there is no edge between $Z$ and $v_{2}$. The undirected edge $v_{1}-v_{2}$ would then be oriented by rule 1 as $v_{1}\to v_{2}$, which leads to a contradiction. Case 2: The two nodes are connected by an undirected edge $Z-v_{2}$. Then, $Z\in T$ and is a parent node of $v_{1}$, contradicting the fact that $v_{1}$ precedes all other nodes in $T$. Case 3: $Z\rightarrow v_{2}$ in $\mathcal{G}$. Meek’s rule 1 would then orient $v_{2}\rightarrow v_{j}$, which would result in an incorrect orientation as $v_{j}$ is assumed to be a parent of $v_{2}$ in $\mathcal{G}_{0}$, again a contradiction. Case 4: $v_{2}\rightarrow Z$ in $\mathcal{G}$. Then the edge $v_{1}-v_{2}$ must be $v_{2}\to v_{1}$ in $\mathcal{G}_{0}$, which contradicts the assumed ordering $v_{1}\prec v_{2}$.
@@
Figure 11: First possibility: orienting $v_{1}\rightarrow v_{j}$ by Line 1 in Algorithm 1 results in contradictions where $v_{1}\rightarrow v_{2}$ is oriented in (a) or $v_{j}\rightarrow v_{2}$ in (b).